# Diffusion Bridge Implicit Models

### v1

@@ D.1 Model Details
For the image inpainting task on ImageNet 256$\times$256 with 128$\times$128 center mask, DDBMs do not provide available checkpoints. Therefore, we train a new model from scratch using the noise schedule of I2SB [25]. The network is initialized from the pretrained class-conditional diffusion model on ImageNet 256$\times$256 provided by [10], while additionally conditioned on $\bm{x}_{T}$. The data prediction model in this case is parameterized by the network $\bm{F}_{\theta}$ as $\bm{x}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})=\bm{x}_{t}-\sigma_{t}\bm{F}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})$ and trained by minimizing the loss $\mathcal{L}(\theta)=\mathbb{E}_{t,\bm{x}_{0},\bm{x}_{T}}\left[\frac{1}{\sigma_{t}^{2}}\|\bm{x}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})-\bm{x}_{0}\|_{2}^{2}\right]$. We train the model on 8 NVIDIA A800 GPU cards with a batch size of 256 for 400k iterations, which takes around 19 days.
D.2 Sampling Details
We elaborate on the sampling configurations of different approaches, including the choice of timesteps $\{t_{i}\}_{i=0}^{N}$ and details of the samplers. In this work, we adopt $t_{\min}=0.0001$ and $t_{\max}=1$ following [53]. For the DDBM baseline, we use the hybrid, high-order Heun sampler proposed in their work with an Euler step ratio of 0.33, which is the best performing configuration for the imgae-to-image translation task. We use the same timesteps distributed according to EDM [21]’s scheduling $(t_{\max}^{1/\rho}+\frac{i}{N}(t_{\min}^{1/\rho}-t_{\max}^{1/\rho}))^{\rho}$, consistent with the official implementation of DDBM. For DBIM, since the initial sampling step is distinctly forced to be stochastic, we specifically set it to transition from $t_{\max}$ to $t_{\max}-0.0001$, and employ a simple uniformly distributed timestep scheme in $[t_{\min},t_{\max}-0.0001)$ for the remaining timesteps, across all settings. For interpolation experiments, to enhance diversity, we increase the step size of the first step from 0.0001 to 0.01.
D.3 License
@@
DIODE-Outdoor https://diode-dataset.org/ [46] MIT
@@
Guided-Diffusion https://github.com/openai/guided-diffusion [10] MIT
I2SB https://github.com/NVlabs/I2SB [25] CC-BY-NC-SA-4.0
@@
Appendix E Additional Samples

### v2

@@ F.1 Model Details
For the image inpainting task on ImageNet 256$\times$256 with 128$\times$128 center mask, DDBMs do not provide available checkpoints. Therefore, we train a new model from scratch using the noise schedule of I2SB (Liu et al., 2023). The network is initialized from the pretrained class-conditional diffusion model on ImageNet 256$\times$256 provided by (Dhariwal & Nichol, 2021), while additionally conditioned on $\bm{x}_{T}$. The data prediction model in this case is parameterized by the network $\bm{F}_{\theta}$ as $\bm{x}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})=\bm{x}_{t}-\sigma_{t}\bm{F}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})$ and trained by minimizing the loss $\mathcal{L}(\theta)=\mathbb{E}_{t,\bm{x}_{0},\bm{x}_{T}}\left[\frac{1}{\sigma_{t}^{2}}\|\bm{x}_{\theta}(\bm{x}_{t},t,\bm{x}_{T})-\bm{x}_{0}\|_{2}^{2}\right]$. We train the model on 8 NVIDIA A800 GPU cards with a batch size of 256 for 400k iterations, which takes around 19 days.
F.2 Sampling Details
We elaborate on the sampling configurations of different approaches, including the choice of timesteps $\{t_{i}\}_{i=0}^{N}$ and details of the samplers. In this work, we adopt $t_{\min}=0.0001$ and $t_{\max}=1$ following (Zhou et al., 2023). For the DDBM baseline, we use the hybrid, high-order Heun sampler proposed in their work with an Euler step ratio of 0.33, which is the best performing configuration for the image-to-image translation task. We use the same timesteps distributed according to EDM (Karras et al., 2022)’s scheduling $(t_{\max}^{1/\rho}+\frac{i}{N}(t_{\min}^{1/\rho}-t_{\max}^{1/\rho}))^{\rho}$, consistent with the official implementation of DDBM. For DBIM, since the initial sampling step is distinctly forced to be stochastic, we specifically set it to transition from $t_{\max}$ to $t_{\max}-0.0001$, and employ a simple uniformly distributed timestep scheme in $[t_{\min},t_{\max}-0.0001)$ for the remaining timesteps, across all settings. For interpolation experiments, to enhance diversity, we increase the step size of the first step from 0.0001 to 0.01.
F.3 License
@@
DIODE-Outdoor https://diode-dataset.org/ Vasiljevic et al. (2019) MIT
@@
Guided-Diffusion https://github.com/openai/guided-diffusion Dhariwal & Nichol (2021) MIT
I2SB https://github.com/NVlabs/I2SB Liu et al. (2023) CC-BY-NC-SA-4.0
@@
Appendix G Additional Samples